OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected! Answer for Blank 1:

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + (x + 1) + (x + 2 ) = 267 }\) Largest integer would be x + 2, obviously.

Parth (parthkohli):

Consecutive = in a row... so there'd be a number, then a number greater than one, then number greater than 2.

OpenStudy (lgbasallote):

let x be the first number since it's "consecutive" that means the next number would be x+1 then the next consecutive number would obviously be x+2 it says the sum of these is 267 therefore you get (x) + (x+1) + (x+2) = 267 you'll want to know what x is first x + x + 1 + x +2 = 267 3x + 3 = 267 3x = 267 - 3 3x = 264 solve for x

OpenStudy (lgbasallote):

only then will you find the largest number (which is x + 2)

OpenStudy (lgbasallote):

and i assure you a numerical answer WILL be obtained